A shopper in a supermarket pushes a cart with a Q/C force of 35.0 N directed at an angle of 25.08 below the horizontal. The force is just sufficient to balance various friction forces, so the cart moves at constant speed. (a) Find the work done by the shopper on the cart as she moves down a 50.0-m-long aisle. (b)The shopper goes down the next aisle, pushing horizontally and maintaining the same speed as before. If the friction force doesn't change, would the shop-per's applied force be larger, smaller, or the same? (c)What about the work done on the cart by the shopper?
1
Expert's answer
2020-10-12T07:44:55-0400

a)


"W=Fs\cos{25.08}=(35)(50)\cos{25.08}=1590\ J"

b)

If the friction force doesn't change, the shopper's applied force be smaller.


"F'=F\cos{25.08}<F"

c)


"W'=W=1590\ J"

The work done on the cart by the shopper would be the same. It is because  the friction force doesn't change.